# The Percept-V Challenge: Can Multimodal LLMs Crack Simple Perception Problems?

### v2

@@ 4.1 Main Results
This section answers our main experimental question (Q1, Q2 and Q3). Table 3 presents the performance of various LLMs across different skills (aggregated over domains that test that skill), as well as over each individual domain. There are several observations to make.
@@
For a significant majority of the domains, across LLMs the performance is below par. In general, as expected, closed source models do better than open source. Among closed source models, GPT-4o which is a non-thinking model, does the worst, with its performance on more than 12 (out of a total of 30) domains 20% or less, and on 24 domains 50% or less. Thinking models do somewhat better, with these numbers being 2 (out of 30), and 17, respectively, for o4-mini, which is closely followed by Gemini. Open source models are in general below par with 20 domains performing 20% or less, and only a single domain performing better than 50%, for Qwen. DeepSeek has similar performance. In terms of overall average performance across domains as well, o4-mini is the best performing model with its accuracy only at 52.7%.
@@
These results clearly point to the inability of LLMs, both closed source, and open, as well thinking and non-thinking, to be able to achieve any meaningful performance on most of the domains in our Percept-V dataset. This is contrast with humans, which perform significantly better on these tasks (see Section 4.3). We also see that LLMs are relatively consistent in their performance across domains, for example, ’inside circles’ has best overall performance, and this is also the domain, where most LLMs achieve their (near) best. Similarly, ’colors present’ has overall the worst performance, and this is also the domain, where individual LLMs also perform poorly compared to other domains. This domain requires models to count objects of a given color, and we hypothesize that MLLMs get confused in related colors (such as turqouise and blue). Overall, this suggests that some domains are inherently harder than others for all LLMs.
Skill-wise Analysis:
A skill-wise analysis of performance presents no different story. o4-mini is best performing with average accuracy across skills at 52.19% closely followed by Gemini. [^2] We note that average accuracy across skills may not match average across across domains, since the same domain may be repeated across multiple skills. The best performing o4-mini can achieve more than 50% accuracy only on 3 out of a total of 7 skills, with its higher accuracy on any skill being 63%, demonstrating that it struggles on majority of the skills to get a reasonable performance. GPT4o sees a sharp decline in performance with average accuracy at 28.6%, and among open source models, Qwen is only at 14.48%. We see that o4-mini and Gemini are consistently better than GPT4o across skills, followed by open source models, which perform quite poorly. Similar to the case of domains, there is consistent behavior of LLM performance skill-wise, with some skills being harder (easier) than others for all LLMs.
For instance, most LLMs perform their best on Visual Spatial Relationship, and (near) worst on Visual Closure. Visual spatial relationships generally test relations such as an object being to the left of or above another object. Since image captions generally contain such relationships, it is likely that MLLMs got trained well on this skill and hence perform well. On the other hand, visual closure questions test on inside-out relationships, layered objects, or matching a solid object with just its outline. Such phenomena are likely absent from multimodal training data, making models weaker on this skill.
@@ 4.2 Performance vs Size
In this section, we set out to answer Q4: what is the variance in LLM performance across varying size? Results comparing the skill-wise model performance across different problem sizes are presented in Figure 2. There are 7 graphs, one for each skill, and the last graph in the bottom row presents performance of each LLM averaged over skills. We observe that performance of models drop significantly with the increase in problem size for all skills. We observe that for most of the skills, thinking models (o4-mini and Gemini) solve smaller problems with lesser number of objects sufficiently well, with near-perfect accuracy, for several skills. However, this performance drops dramatically with increasing size. As seen in Table 3, o4-mini and Gemini are among the best performing models, and are relatively robust to size change compared to other LLMs, despite a performance drop. These are followed by GPT-4o, and open source models perform the worst. Clearly, these results indicate that LLMs are not able to generalize across size across different skills. We hypothesize that the key reason for this performance drop is their inability to ’discern’ the underlying pattern of the task to be solved in a given domain, resulting in sharp decline in performance as the size increases. It is also possible that for thinking models, they are able to ’discern’ this pattern, but are unable to operationalize this for larger sizes.
@@ 4.3 Human Study
This section answers Q5, i.e., how does LLM performance vary in contrast with human subjects. A study on 16 college students was done with the Percept-V. It had 6 female and 10 male subjects. The subjects were presented randomly selected questions of any problem size from the 30 domains. Results of Human Study are presented in Figure 3. Here, for fair comparison, the models are also evaluated on the exactly same samples as were given to the humans. It can be seen that models lag severely as compared to humans. Even the best performance of the models exhibited by Gemini in Visual Spatial Relationship is around 60% as compared to about 90% accuracy shown by humans in the same skill. Overall see Table 4 – human performance is about 30 percent points higher than than closest MLLM. When testing on ‘colours present’, the domain on which all MLLMs have abysmal performance, we find the gap to be humongous – humans achieve 81.25% performance whereas best MLLMs on the same questions achieve only 6.25%.
This experiment underscores a large gap in the overall perceptual ability of MLLMs compared to humans – an important direction for future work.
Model Overall performance
GPT-4o 31.17%
o4-mini 51.46%
Gemini 49.27%
Qwen 20.40%
Deepseek 8.45%
Human 81.67%
Table 4: Overall performance in human study
4.4 Compliance with Output Format
@@
Answer Type GPT-4o o4-mini Gemini Qwen DeepSeek
single 0.00 0.00 0.00 0.00 3.68
fixed 0.00 0.17 0.83 5.18 61.58
list 0.00 0.00 0.00 0.00 22.44
set 0.00 0.10 0.20 19.10 60.00
@@ 5 Conclusion
Our paper takes an important step forward in assessing the perceptual abilities of multimodal LLMs. To isolate perceptual ability, our questions use very few and basic reasoning skills (such as counting), and similarly very basic knowledge (of colors and basic shapes). We present Percept-V, an extensive dataset of 6000 automatically generated images from 30 domains, which are uncontaminated, i.e., not seen by any LLM before. We also annotate each domain with their perceptual skills from the TVPS-4 Cognitive Science framework.
When tested on our dataset, all MLLMs perform rather weakly on almost all domains. The best average performance on our dataset is only 52.7% (o4-mini), and open models perform rather poorly (8-18%). As complexity of questions increases, all model performances degrade quite fast on almost all skills, with thinking models continuing to perform decently on a couple of skills.
We also perform human study to compare human performance with MLLMs. We find that the gap between the best MLLM and humans is quite high (over 30 percent points), suggesting that contrary to LLMs’ high performance against humans in many tasks (e.g., LLMs winning Math olympiad), MLLMs are rather behind, especially on basic perception skills. We will release Percept-V upon publication. We hope that it will become a standard dataset for testing any new MLLM, and will help in improving the quality of future MLLMs.
@@ A.1 Domain examples of MLLMs
A.1.3 circle_location
Prompt:
@@
A.1.4 circle_right_triangle
Prompt:
@@
A.1.5 colours_present
Prompt:
@@ A.1.6 Comparing Size
- This is an image with several rows of circles where each row is above a horizontal line.
- Each row contains two colored circles (one blue and one green) of different sizes.
- Analyse which circle is bigger between the two in each row.
- The last line of the output should be of form "ANSWER:" followed by the space separated color answers for all the rows in the image.
- For example ANSWER: Blue Green Blue.
@@ A.1.7 Count Colored Circles
- This is an image containing some colored circles.
- This is an image containing some colored circles.
- Last line of the output must be of the form COUNT:x
@@ A.1.8 Counting Circles
- This is an image containing some black circles.
- Count the total number of circles in the image.
- The last line of the output should be of form COUNT:x.
@@ A.1.9 Counting Locations
- There will be an image containing a rectangle and some circles above and below it.
- Count the number of circles above the rectangle and below the rectangle.
- The output must be of form in a single line ABOVE:x BELOW:y
@@ A.1.10 Counting Shapes
- This is an image containing shapes such as circles, triangles and squares.
- Count the number of circles, triangles and sqaures in the image.
- The last line of the output must be of the form CIRCLES:x TRIANGLES:y SQUARES:z
@@ A.1.11 Cross and Knots
- This is a grid containing some black X.
- Each cell is given a coordinate (x,y) which is (row,column) with 0 based indexing.
- An adjacent cell to a cell A is any cell sharing an edge with A
- A safe cell is a cell that is not adjacent to any black X.
- Output the coordinates of any one safe cell as labelled in the grid.
- If no such cell exists output None.
- The last line of the output should just contain the cell coordinates (x,y) or None and nothing else.
@@ A.1.12 Graph Counting
- This is a graph where the blue colored circles are nodes and the lines joining them are edges.
- Count the number of nodes and edges in the graph.
- Last line of the output must be of the form NODES:x EDGES:y.
@@ A.1.13 Grid Path
- This is a grid which contains some circles, triangles and squares.
- Travel from cell labelled S to cell labelled E.
- Follow the path marked by the black line and arrows.
- Write down the sequence of shapes seen while moving along this path.
- The last line of the output should be a comma-separated list of shapes without colors, written in order of visiting of form SHAPES : Shape 1 , Shape 2 and so on.
@@ A.1.14 Identifying Shapes
- This is an image with several rows of different shapes where each row is above a horizontal line. - Each row contains a coloured shape.
- Identify the shape present in each row.
- The ouput should have only a single line containing the answer.
- The output format should be "ANSWER:" followed by the space seperated shape names.
- For example, ANSWER: Circle Triangle Circle
@@ A.1.15 Inside Circles
- This is an image containing some black circles and a single red dot.
- Count the total number of circles in the image which contain the red dot
- Note that a circle containing the red dot means the centre of the dot is present inside the circumference of the circle
- The last line of the output should be of form COUNT:x.
@@ A.1.16 Layered Colors
- This image contains some concentric circles in different colours.
- Write down the sequence of colours seen from inside to outside.
- The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia]
- No colour is repeated
- The output should be written in a single line as a comma-separated list of colours, written in order from the innermost colour to the outermost colour
- For example, the output should be of the form COLOURS: Colour 1 , Colour 2 and so on.
@@ A.1.17 Layered Shapes
- This image contains some shapes that are layered on top of each other in a white background.
- The image could also contain a single shape with no other shapes layered on top of it.
- Write down the sequence of shapes from the inside layer to the outside layer.
- If there is a single shape then identify it.
- The last line of the output should be a comma-separated list of shapes without colors, written in order from top to bottom of form SHAPES : Shape 1 , Shape 2 and so on.
- If there is a single shape then write SHAPES : shape-name where shape-name is the name of the shape seen
@@ A.1.18 List Colors
- This is an image with some multi-coloured shapes drawn on top of it
- The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia]
- No colour is repeated
- Identify the colour of the shapes drawn over the background image
- Form a list of all the colours of the shapes
- The output must be given in a single line in the form of a list of all the colours found
@@ A.1.19 List Shapes
- This is an image with some multi-coloured shapes drawn on top of it.
- Identify the shapes drawn over the background image
- Form a set of all the unique shapes seen in the image
- The output must be given in a single line in the form of a set of unique shapes found.
@@ A.1.20 Locate Circles Color
- This is a grid which contains some coloured circles.
- The rows and columns are numbered in the grid.
- The coordinate of a cell is given by the row number followed by the column number.
- Write down all the coordinates of the cells which contain a green circle.
- The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle.
- The coordinates should be in the format (row,column) withh the comma and bracket.

### v3

@@ 4.1 Main Results
This section answers our main research questions (Q1, Q2 and Q3). Table 3 presents the performance of various LLMs across different skills (aggregated over domains that test that skill), as well as over each individual domain.
@@
For a significant majority of the domains, across MLLMs the performance is below par. In general, as expected, closed source models do better than open source. Among closed source models, GPT-4o which is a non-thinking model, does the worst, with its performance on more than 12 out of 30 domains is 20% or less, and on 24 domains 50% or less. Thinking models do somewhat better, with these numbers being 2 and 17, respectively, for o4-mini, which is closely followed by Gemini and GPT-5-mini. Open source models are in general below par with 20 domains obtaining 20% or less accuracy, and only a single domain performing better than 50%, for Qwen. DeepSeek has similar performance. In terms of overall average performance across domains as well, GPT-5-mini is the best performing model, but with its accuracy only at 55.22%.
These results clearly point to the inability of MLLMs, both closed source, and open, as well thinking and non-thinking, in achieving any meaningful performance on most of the domains in our Percept-V dataset. This is in contrast with humans, who perform significantly better on these tasks (see Section 4.4). We also see that MLLMs are relatively consistent in their performance across domains, for example, ‘inside circles’ has best overall performance, and this is also the domain, where most MLLMs achieve their (near) best. Similarly, ‘colors present’ has overall the worst performance, and this is also the domain, where individual MLLMs also perform poorly compared to other domains. This domain requires models to count objects of a given color, and we hypothesize that MLLMs get confused in related colors, such as turqouise and blue (even though a legend clarifies all colors at the top of each image). Overall, this suggests that some domains are inherently harder than others for all current-day MLLMs.
Skill-wise Analysis:
A skill-wise analysis of performance presents no different story. GPT-5-mini performs the best, with average accuracy across skills at 54.77%, closely followed by o4-mini and Gemini. [^2] A domain is included in all skills it exhibits GPT-5-mini achieves more than 50% accuracy only on 4 out of 7 skills, with its higher accuracy on any skill being 66.64%, demonstrating that it struggles to get a reasonable performance on all skills. GPT-4o sees a sharp decline in performance with average accuracy at 28.6%, and among open source models, Qwen is only at 14.48%. We see that o4-mini and Gemini are consistently better than GPT-4o across skills, followed by open source models, which perform quite poorly. Similar to domain-wise analysis, there is consistent behavior of MLLM performance skill-wise, with some skills being harder (easier) than others for all MLLMs.
For instance, most LLMs perform their best on Visual Spatial Relationship, and (near) worst on Visual Closure. Visual spatial relationships test relations such as an object being to the left of or above another object. Since image captions generally contain such relationships, it is likely that MLLMs got trained well on this skill. On the other hand, visual closure questions test on inside-out relationships, layered objects, or matching a solid object with just its outline. Such phenomena are likely absent from multimodal training data, making models weaker on this skill.
@@ 4.2 Performance vs Size
We now answer Q4: what is the variance in LLM performance across varying size? Results comparing the skill-wise model performance across different problem sizes are presented in Figure 2. There are 7 graphs, one for each skill, and the last graph in the bottom row presents performance of each MLLM averaged over skills. We observe that performance of models drop significantly with the increase in problem size for all skills. We observe that for most of the skills, thinking models (o4-mini and Gemini) solve smaller problems with less number of objects sufficiently well, with near-perfect accuracy, for several skills. However, this performance drops dramatically with increasing size. As seen in Table 3, GPT-5-mini, o4-mini and Gemini are among the best performing models, and are relatively robust to size change compared to other LLMs, despite a performance drop. These are followed by GPT-4o, and open source models perform the worst. Clearly, these results indicate that LLMs are not able to generalize across sizes very well.
4.3 Alternate Skill Analysis
Separate from TVPS-4 classification, we observe that our dataset uses three other frequent skills – counting, optical character recognition (OCR) and handling grid-like images. They appear in 13, 9, and 7 domains, respectively. We note that counting is a basic reasoning skill, employed primarily so as to ease automatic evaluation of answers, nevertheless, it may distract from the main results. An analysis of counting (C) vs non-counting (NC) domains shows the following results: GPT-4o (C:29.03, NC:32.13), GPT5-mini (C: 47.73, NC: 58.54), o4-mini (C:42.99, NC:58.44), Gemini (C:40.92, NC:55.22), Qwen (C:19.88, NC:15.93), and DeepSeek (C:5.65, NC:6.88). This suggests that while performance is not very dissimilar, for some thinking MLLMs, inclusion of counting skill may make the problem somewhat harder. Still, even for non-counting domains, the performance is not particularly stronger, so it is safe to conclude counting ability may be one of the factors, but does not explain all of the weak results on our dataset.
Figure 3: The overall accuracy of all models in counting, grid understanding and ocr skills.
Figure 3 shows performance for the three alternate skills vs. problem size. We observe that models have a similar sharp drop in case of counting and OCR. However, the case for grid understanding is different as models, particularly, reasoning models like GPT-5-mini, o4-mini and Gemini have a more-or-less plateau-like curve with no significant accuracy drops. In fact, GPT-5-mini shows 100% accuracy in 3 out of 7 grid-based domains! We hypothesize that the training data distribution (especially in RLVR tasks) may frequently include grid-like structures, and that might have resulted in their better understanding of grids.
@@ 4.4 Human Study
We now answer Q5 – how do MLLMs compare with human performance. We recruit 16 college students, 6 female and 10 male. They were presented randomly selected questions of any problem size from the 30 domains. Results of this study are presented in Figure 4. For fair comparison, the models are also evaluated on the exactly same samples, and exactly the same protocol, as the humans. It can be seen that models lag severely as compared to humans. Even the best performance of the models exhibited by Gemini in Visual Spatial Relationship is around 73.08% as compared to about 96.42% accuracy shown by humans in the same skill. Overall (see Table 5 in appendix) – human performance is more than 40% points higher than the closest MLLM. Interestingly, when testing on ’colours_present’, the domain on which all MLLMs have abysmal performance, we find the gap to be humongous – humans achieve 90% performance whereas best MLLMs on the same questions achieve only 6.67%.
4.5 Other Results
@@
Answer Type GPT-5-mini GPT-4o o4-mini Gemini Qwen DeepSeek
single 0.14 0.00 0.00 0.00 0.00 3.68
fixed 0.00 0.00 0.17 0.83 5.18 61.58
list 2.19 0.00 0.00 0.00 0.00 22.44
set 0.00 0.00 0.10 0.20 19.10 60.00
@@
One-shot Results:
We additionally perform one-shot experiments on proprietary models to assess how much exemplars can help MLLMs in understanding the task better. Contrary to conventional wisdom, we find that one-hot prompting degrades the overall performance, with average of skills accuracy reducing to 51.25% for GPT-5-mini, 17.62% for GPT-4o, 52.26% for o4-mini and 44.42% for Gemini (refer to Table 6 in Appendix). Moreover, the skill-wise graphs also show degrading trends across sizes, similar to zero-shot (see Fig. 35 in Appendix). This hints that MLLMs’ weak performance is not likely because of task understanding, but due to inherent ability limitations.
@@ 5 Conclusion
Our paper takes an important step forward in assessing the perceptual abilities of multimodal LLMs. We present Percept-V, an extensive dataset of 6000 automatically generated images from 30 domains where each domain is mapped with a subset of perceptual skills from the TVPS-4 Cognitive Science framework. To isolate perceptual ability, questions in our dataset makes use very few basic reasoning skills (such as counting), and similarly very basic knowledge (of colors and basic shapes). A variety of open and closed source MLLMs perform rather weakly on almost all domains, with best average performance of any model being 55.2%, demonstrating significant gap in their ability to solve basic perceptual tasks. Our human study experiments show that the gap between the best MLLM and humans is quite high suggesting that contrary to LLMs’ high performance against humans in existing tasks (e.g., LLMs winning Math olympiad), MLLMs are rather behind, especially on basic perception skills. We will release Percept-V upon publication, and hope that it will become a de-facto standard for benchmarking perceptual abilities of any new and existing MLLMs.
@@ A.1.1 Change Color
Figure 5: Change Color
Figure 6: Change Color
@@ A.1.2 Circle Boxes
Figure 7: Circle Boxes
A.1.3 Circle Location
Input Prompt:
@@ A.1.3 Circle Location
Figure 8: Circle Location
A.1.4 Circle Right Triangle
Input Prompt:
@@ A.1.4 Circle Right Triangle
Figure 9: Circle Right Triangle
A.1.5 Colours Present
Input Prompt:
@@ A.1.6 Comparing Size
- This is an image with several rows of circles where each row is above a horizontal line. - The image may also contain a single row with a blue and a green circle. - Each row contains two colored circles (one blue and one green) of different sizes. - Analyse which circle is bigger between the two in each row. - The last line of the output should be of form "ANSWER:" followed by the space separated color answers for all the rows in the image. - For example ANSWER: Blue Green Blue
@@ A.1.7 Count Colored Circles
- This is an image containing some colored circles - Count the total number of red circles in the image - Last line of the output must be of the form COUNT:x
Figure 11: Count Colored Circles
@@ A.1.8 Counting Circles
- This is an image containing some black circles - Count the total number of circles in the image - The last line of the output should be of form COUNT:x.
Figure 12: Counting Circles
@@ A.1.9 Counting Locations
- There will be an image containing a rectangle and some circles above and below it - Count the number of circles above the rectangle and below the rectangle - The output must be of form in a single line ABOVE:x BELOW:y
Figure 13: Counting Locations
@@ A.1.10 Counting Shapes
- This is an image containing shapes such as circles, triangles and squares - Count the number of circles, triangles and sqaures in the image - The last line of the output must be of the form CIRCLES:x TRIANGLES:y SQUARES:z
Figure 14: Counting Shapes
@@ A.1.11 Cross and Knots
- This is a grid containing some black X. - Each cell is given a coordinate (x,y) which is (row,column) with 0 based indexing. - An adjacent cell to a cell A is any cell sharing an edge with A - A safe cell is a cell that is not adjacent to any black X. - Output the coordinates of any one safe cell as labelled in the grid. - If no such cell exists output None. - The last line of the output should just contain the cell coordinates (x,y) or None and nothing else.
Figure 15: Cross and Knots
@@ A.1.12 Graph Counting
- This is a graph where the blue colored circles are nodes and the lines joining them are edges. - The graph may contain a single blue node with no edges. - The graph may be disconnected. - Count the number of nodes and edges in the graph. - Last line of the output must be of the form NODES:x EDGES:y.
Figure 16: Graph Counting
@@ A.1.13 Grid Path
- This is a grid which contains some circles, triangles and squares. - Travel from cell labelled S to cell labelled E. - Follow the path marked by the black line and arrows. - Write down the sequence of shapes seen while moving along this path. - The shapes present in the S and E cells should also be included in the sequence. - The last line of the output should be a comma-separated list of shapes without colors, written in order of visiting of form SHAPES : Shape 1 , Shape 2 and so on.
Figure 17: Grid Path
@@ A.1.14 Identifying Shapes
- This is an image with several rows of different shapes where each row is above a horizontal line. - Each row contains a coloured shape. - The image may also contain a single shape.
- Identify the shape present in each row. - In case there is a single shape, treat it as a single row. - The ouput should have only a single line containing the answer. - The output format should be "ANSWER:" followed by the space seperated shape names. - For example, ANSWER: Circle Triangle Circle
@@ A.1.15 Inside Circles
- This is an image containing some black circles and a single red dot - Determine if the red dot is contained inside any circle in the image - Note that a circle containing the red dot means the centre of the dot is present inside the circumference of the circle - The last line of the output should be of form ANSWER: x where x is Yes if the dot is contained in any circle and No otherwise.
Figure 18: Inside Circles
@@ A.1.16 Layered Colors
- This image contains some concentric circles in different colours. - The image may also contain a single circle - Write down the sequence of colours seen from inside to outside. - In case of a single circle in the image, write down the colour of that circle. - The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia] - No colour is repeated - The output should be written in a single line as a comma-separated list of colours, written in order from the innermost colour to the outermost colour - For example, the output should be of the form COLOURS: Colour 1 , Colour 2 and so on.
Figure 19: Layered Colors
@@ A.1.17 Layered Shapes
- This image contains some shapes that are layered on top of each other in a white background. - The shapes colud be of the following types - [circle, diamond, hexagon, octagon] - The image could also contain a single shape with no other shapes layered on top of it. - Write down the sequence of shapes from the inside layer to the outside layer. - If there is a single shape then identify it. - The last line of the output should be a comma-separated list of shapes without colors, written in order from top to bottom of form SHAPES : Shape 1 , Shape 2 and so on. - If there is a single shape then write SHAPES : shape-name where shape-name is the name of the shape seen
Figure 20: Layered Shapes
@@ A.1.18 List Colors
- This is an image with some multi-coloured shapes drawn on top of it - The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia] - No colour is repeated - Identify the colour of the shapes drawn over the background image - Form a list of all the colours of the shapes - The output must be given in a single line in the form of a list of all the colours found
Figure 21: List Colors
@@ A.1.19 List Shapes
- This is an image with some multi-coloured shapes drawn on top of it - The shapes belong to the following list - [circle, triangle, square, pentagon] - Identify the shapes drawn over the background image - Form a set of all the unique shapes seen in the image - The output must be given in a single line in the form of a set of unique shapes found
Figure 22: List Shapes
@@ A.1.20 Locate Circles Color
- This is a grid which contains some coloured circles. - The rows and columns are numbered in the grid. - The coordinate of a cell is given by the row number followed by the column number. - Write down all the coordinates of the cells which contain a green circle. - The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle. - The coordinates should be in the format (row,column) with the comma and bracket
Figure 23: Locate Circles Color